# Hamiltonian Active Matter in Incompressible Fluid Membranes

### v3

@@ Abstract
Active proteins and membrane-bound motors exert force dipole flows along fluid interfaces and lipid bilayers. We develop a unified hydrodynamic and Hamiltonian framework for the interactions of pusher and puller dipoles embedded in an incompressible two-dimensional membrane supported by a shallow viscous subphase. Beginning from the screened Stokes equations of the membrane–subphase composite, we derive the real-space incompressible Green’s tensor, obtain its near- and far-field asymptotics, and construct the resulting dipolar velocity and stream functions. Although generic dipoles reorient under the local membrane vorticity, we show that the far-field dipolar flow is vorticity-free; force-free motors therefore retain fixed orientation and obey a Hamiltonian dynamics in which the positions of $N$ dipoles evolve via an effective Hamiltonian built from the dipolar stream function. In the near field, where the flow possesses finite vorticity, a Hamiltonian formulation is recovered in the quenched-orientation limit. Exploiting this structure, we simulate ensembles of pusher and puller dipoles and compare the dynamics generated by the $1/r$ near-field kernel and the subphase screened $1/r^{3}$ far-field kernel. For identical dipoles, the far-field Hamiltonian produces rapid clustering from random initial conditions, whereas the near-field Hamiltonian suppresses collapse and yields extended, non-aggregating configurations.
@@ I Introduction
Supported membranes provide a particularly interesting setting in which these hydrodynamic effects have been worked out in detail. Classical analyses of membrane drag by Saffman, Delbrück, and their successors [11, 12, 13, 14, 15, 16, 17] showed that flows in a viscous membrane coupled to a finite–depth subphase obey a screened Stokes equation. The resulting Green’s function exhibits two qualitatively distinct regimes: (i) a logarithmic “near field” with unscreened in–plane hydrodynamics producing $1/r$ dipolar flows, and (ii) a “far field” in which substrate friction produces algebraically screened $1/r^{3}$ flows. Recent studies of swimmers and active inclusions in thin films and membranes [18, 19, 20, 21] highlight the richness of these interactions, but a systematic comparison of near– and far–field dipolar dynamics in supported membranes has been lacking.
In this work we develop a unified hydrodynamic and Hamiltonian description of pusher and puller interactions in an incompressible supported membrane. Beginning from the screened Stokes equations of the membrane–subphase composite, we obtain the real–space Green’s tensor, derive its near– and far–field asymptotics, and construct the associated dipolar velocity and stream functions. Although a generic dipole reorients in response to the local membrane vorticity, we show that the far–field dipolar flow is vorticity free. Force–free motors therefore maintain fixed orientation and admit a reduced Hamiltonian description closely related to recent two–dimensional Hamiltonian formulations of active particles [22]. In the near field, where vorticity is finite, a Hamiltonian structure is recovered in the quenched–orientation limit.
Exploiting this framework, we compare the collective behaviour generated by the unscreened and screened kernels for ensembles of identical pushers and pullers. The contrast is striking: the far–field Hamiltonian produces rapid collapse into compact clusters, whereas the near–field Hamiltonian suppresses aggregation and yields extended, non–collapsing configurations. These trends persist across a wide range of initial geometries and motor types, demonstrating that the structure of the hydrodynamic kernel—not the sign of the dipole alone—determines whether a membrane–bound active ensemble aggregates or disperses.
Taken together, our results show that even in the simplest incompressible limit, the hydrodynamic screening intrinsic to supported membranes imposes strong, kernel–dependent constraints on collective organisation. This work provides a foundation for Hamiltonian approaches to active matter on interfaces and sets the stage for extensions to chiral membranes, curved geometries, and biologically driven active films [20, 21, 23, 24].
@@ II Incompressible supported membrane
We consider a two–dimensional viscous membrane of shear viscosity $\eta_{s}$ lying in the plane $z=h$, supported by a Newtonian subphase of viscosity $\eta$ and thickness $h$ above a rigid wall (Fig. 1). Throughout the main text we restrict to the incompressible limit: the in–plane velocity field $\mathbf{v}(\mathbf{r})$ satisfies $\nabla\!\cdot\!\mathbf{v}=0$, there is no dilatational viscosity.
@@
In Fourier space, incompressibility projects onto transverse components, and the Brinkman operator becomes
@@
where $K_{n}$ are modified Bessel functions of the second kind. Equations (7)–(9) constitute the screened two–dimensional Stokeslet for an incompressible supported membrane [16, 17].
@@ III Dipolar flows: near and far zones
III.1 Near field: unscreened logarithmic hydrodynamics
@@
characteristic of a two–dimensional stresslet with logarithmic hydrodynamics.
@@
III.2 Far field: algebraically screened potential flow
@@
which smoothly matches the near–field expression at $r\sim\kappa^{-1}$. In contrast to the near zone, the far–field flow is irrotational away from the dipole singularity:
@@ IV.1 Near–field Hamiltonian
As noted in Eq. (15), the unscreened near–field dipolar flow carries finite vorticity, so freely rotating dipoles undergo hydrodynamic reorientation. These orientational dynamics preclude a Hamiltonian written solely in terms of particle positions. A reduced Hamiltonian description is nevertheless recovered in the quenched–orientation limit, where dipole axes are externally constrained or relax much faster than translational motion. This regime is natural for anchored “shakers” and for strongly aligned active suspensions [3, 4, 5, 22]. With orientations fixed, the near–field stresslet retains its $1/r$ form and generates a consistent position-based Hamiltonian, enabling a direct comparison with the far–field dynamics. Thus, in the unscreened regime ($\kappa r_{ij}\ll 1$), the stream function of dipole $j$ is given by Eq. (14),
$\Psi_{j}^{\rm(near)}=\frac{\sigma_{j}}{8\pi\eta_{s}}\sin\!\big[2(\theta_{ij}-\alpha_{j})\big],$ (22)
leading to the Hamiltonian
@@ V Collective dynamics of pusher and puller dipoles
We probe the dynamical consequences of the two kernels by separately simulating ensembles of identical pushers and pullers ($\sigma_{i}=\pm 1$) with fixed, co–aligned orientations. In the far field, the flow is vorticity–free and orientations are naturally quenched, making $H_{\rm far}$ an exact position–based Hamiltonian. In the near field, where vorticity is finite, a Hamiltonian description applies only in the enforced quenched–orientation limit. Within these regimes, we contrast the collective organization produced by $H_{\rm far}$ and $H_{\rm near}$.
@@ V.1 Simulation details
The collective dynamics of active dipoles in an incompressible membrane depend crucially on whether interactions are mediated by the algebraically screened far–field kernel or by the unscreened near–field kernel. Figure 2 summarizes the behaviour for ensembles of $N=15$ pushers (top row) and pullers (bottom row).
@@
In contrast, evolution under the unscreened near–field Hamiltonian $H_{\rm near}$ [Figs. 2(c,d,g,h)] exhibits a completely diffrent trend. The $1/r$ stresslet decay is slower and we find that the ensemble does not collapse. Instead, particles spread into extended configurations and $\langle d_{ij}\rangle(t)$ grows monotonically. In this near–zone regime, the self–aggregation observed in the far field is vastly suppressed: although individual pairs can transiently attract or repel, many–body constraints prevent the formation of a stable compact cluster.
The similarity between the pusher and puller panels highlights a central result of this work: the nature of the hydrodynamic kernel, rather than the motor type alone, determines whether active dipoles aggregate or disperse. The far–field screened flow is irrotational and generates an attractive effective Hamiltonian, while the near–field unscreened flow frustrates collective collapse. For completeness, Fig. 3 shows a representative far–zone evolution with arbitrary orientations, illustrating that the same aggregation mechanism persists beyond the co–aligned limit. Representative far- and near-zone trajectories of co-aligned pusher and puller clusters are provided in the Supplementary Material (Figs. S1–S2). In all these examples, a soft harmonic repulsion is added to regularize close encounters and suppress particle overlap.
@@ VI Summary and outlook
We have developed a minimal hydrodynamic Hamiltonian framework for active force dipoles confined to an incompressible supported membrane. Starting from the Brinkman–regularized Stokes equation, we derived the screened incompressible Green’s tensor, identified its logarithmic near–field and algebraically screened far–field limits, and constructed the associated dipolar flows and stream functions. A key outcome of this analysis is the emergence of two distinct Hamiltonian regimes. In the far field, the screened flow is strictly vorticity–free, so dipole orientations remain naturally quenched; membrane incompressibility then yields an exact position–based Hamiltonian $H_{\rm far}$. In contrast, the near field possesses finite vorticity, precluding an orientation–dynamics–free description; a Hamiltonian formulation applies only in the enforced quenched–orientation limit, defining the effective Hamiltonian $H_{\rm near}$. Within these regimes, we contrasted the collective organization generated by the two kernels.
@@
This comparison reveals a striking dichotomy. For identical pushers (and similarly for pullers), the algebraically screened $1/r^{3}$ far–field interactions produce a strong self–binding instability, driving rapid collapse into compact clusters and a sharp decay of the mean pair separation. The unscreened $1/r$ near–field kernel behaves qualitatively differently: it induces strong local rearrangements but no long–lived aggregates, and the ensemble disperses monotonically in time, rendering near–zone dynamics intrinsically non–aggregating. This contrast persists across motor types and initial orientations, showing that it is the hydrodynamic kernel—not merely the dipole sign—that controls collective organization. These results offer a controlled framework for interpreting more complex membrane rheologies. Allowing membrane compressibility opens a longitudinal mobility channel that breaks the Hamiltonian structure and can enhance aggregation for pullers [24] while producing mixed responses for pushers. Introducing odd (Hall) viscosity adds antisymmetric stresses that drive chiral transport and circulating cluster states [24]. Coupling the membrane to curvature, shape fluctuations, or biochemical activity further broadens the phenomenology, enabling a range of richer and biologically relevant collective behaviours. More broadly, the incompressible supported membrane studied here provides a simple yet dynamically rich framework for future works on compressible, chiral, or curved membranes, and it motivates experimental tests in reconstituted active films where the screening length and membrane viscosities can be precisely tuned. The Hamiltonian formulation developed in this work offers a minimal yet powerful analytical framework for probing active matter at interfaces, and highlights the central role of hydrodynamic screening in shaping collective organization.
@@ Acknowledgements.
The data that support the findings of this study are available within the article.
Supplemental Material for
“Hamiltonian Active Matter in Incompressible Fluid Membranes”
(a) Pushers.
(b) Pullers.
Figure S1: Far-zone dynamics of co-aligned clusters of twelve pushers (left) and pullers (right) in an incompressible membrane, initialized at random locations within a disc. Initial positions appear in green (pushers) and yellow (pullers), final positions in red and orange, with trajectories in grey. A soft harmonic repulsion is included to prevent particle overlap.
(a) Pushers.
(b) Pullers.
Figure S2: Representative near-zone dynamics of randomly initialized (within a disc), co-aligned clusters of twelve pushers (left) and pullers (right) in an incompressible membrane. Initial positions are shown in green (pushers) and yellow (pullers); final positions in red and orange, respectively; trajectories are grey. A soft harmonic repulsion is included to regularize close encounters.

### v4

# Hamiltonian Active Particles in Incompressible Fluid Membranes

@@ Abstract
Active proteins and membrane-bound motors exert force dipole flows along fluid interfaces and lipid bilayers. We develop a Hamiltonian framework for the interactions of pusher and puller dipoles embedded in an incompressible two-dimensional membrane supported by a shallow viscous subphase. Beginning from the Brinkman-regularized Stokes equations of the membrane–subphase system, we construct the near- and far-field dipolar velocity and associated stream functions. Although generic dipoles reorient under the local membrane vorticity, we show that the far-field dipolar flow is vorticity-free; force-free motors therefore retain fixed orientation and obey a position-based Hamiltonian dynamics in which the positions of $N$ dipoles evolve via an effective Hamiltonian built from the dipolar stream function. In the near field, where the flow possesses finite vorticity, a Hamiltonian formulation is recovered in the quenched-orientation limit. We demonstrate this Hamiltonian structure by simulating ensembles of pusher and puller dipoles and compare the dynamics generated by the near-field and far-field Hamiltonians. For identical dipoles, the far-field Hamiltonian produces rapid clustering from random initial conditions, whereas the near-field Hamiltonian suppresses collapse and yields extended, non-aggregating configurations. Our work thus provides a concrete realization of position-based Hamiltonian descriptions for active particles in incompressible fluid membranes and establishes the important roles of membrane incompressibility and fluid vorticity that leads to the emergence of such Hamiltonian descriptions.
@@ I Introduction
Supported membranes provide a particularly interesting setting in which these hydrodynamic effects have been worked out in detail. Classical analyses of membrane drag by Saffman, Delbrück, and their successors [11, 12, 13, 14, 15, 16, 17] showed that flows in a viscous membrane coupled to a shallow subphase obey a Brinkman–regularized Stokes equation. The resulting Green’s function exhibits two qualitatively distinct regimes: (i) a logarithmic “near field” with unscreened in–plane hydrodynamics producing $1/r$ dipolar flows, and (ii) a “far field” in which substrate friction produces $1/r^{3}$ dipolar flows. Recent studies of swimmers and active inclusions in thin films and membranes [18, 19, 20, 21] highlight the richness of these interactions, but a concrete realization of a Hamiltonian framework for pusher and puller motors in such incompressible membranes and importantly, the role of membrane fluid vorticity, remains to be explored and clarified. The main objective of this work is to explicitly demonstrate how such a position-based Hamiltonian description naturally emerges for active particles in incompressible fluid membranes.
In this work we develop a Hamiltonian description of pusher and puller interactions in an incompressible membrane coupled to a shallow subphase. Beginning from the screened Stokes equations of the membrane–subphase system, we obtain the real–space Green’s tensor, derive its near– and far–field asymptotics, and construct the associated dipolar velocity and stream functions. Although a generic dipole reorients in response to the local membrane vorticity, we show that the far–field dipolar flow is vorticity free. Force–free motors therefore maintain fixed orientation and admit a reduced position-based Hamiltonian description. However, in the near field, where vorticity is finite, a Hamiltonian structure is recovered in the quenched–orientation limit, closely related to recent two–dimensional Hamiltonian formulations of active particles [22].
Exploiting the Hamiltonian framework, we compare the collective behaviour generated by the near-field and far-field Hamiltonians for ensembles of identical pushers and pullers. The contrast is striking: the far–field Hamiltonian produces rapid collapse into compact clusters, whereas the near–field Hamiltonian suppresses aggregation and yields extended, non–collapsing configurations. These trends persist across a wide range of initial geometries and motor types, demonstrating that the structure of the hydrodynamic kernel—not the sign of the dipole alone—determines whether a membrane–bound active ensemble aggregates or disperses.
Taken together, our results show that even in the simplest incompressible limit, the hydrodynamic screening intrinsic to supported membranes imposes strong, kernel–dependent constraints on collective organisation. This work provides a foundation for similar Hamiltonian approaches to active matter on interfaces and sets the stage for extensions to chiral membranes, curved geometries, and biologically driven active films [20, 21, 23, 24].
@@ II Incompressible supported membrane
We consider a two–dimensional viscous membrane of shear viscosity $\eta_{s}$ lying in the plane $z=h$, supported by a Newtonian shallow subphase of viscosity $\eta$ and thickness $h$ above a rigid wall (Fig. 1). Throughout the main text we restrict to the incompressible limit: the in–plane velocity field $\mathbf{v}(\mathbf{r})$ satisfies $\nabla\!\cdot\!\mathbf{v}=0$, there is no dilatational viscosity.
@@
In Fourier space, membrane incompressibility can be utilized to obtain
@@
where $K_{n}$ are modified Bessel functions of the second kind. Equations (7)–(9) constitute the two–dimensional Stokeslet for an incompressible supported membrane [16, 17].
@@ III Dipolar flows: near and far zones
III.1 Near field: unscreened hydrodynamics
@@
characteristic of a two–dimensional stresslet with logarithmic Stokeslet flow.
@@
III.2 Far field: screened hydrodynamics
@@
which matches the near–field expression at $r\sim\kappa^{-1}$. In contrast to the near zone, the far–field flow is irrotational away from the dipole singularity:
@@ IV.1 Near–field Hamiltonian
As noted in Eq. (15), the unscreened near–field dipolar flow carries finite vorticity, so freely rotating dipoles undergo hydrodynamic reorientation. These orientational dynamics preclude a Hamiltonian written solely in terms of particle positions. A reduced Hamiltonian description is nevertheless recovered in the quenched–orientation limit, where dipole axes are externally constrained or there is a separation in orientation relaxation timescales compared to translational motion. This regime is natural for anchored “shakers” and for strongly aligned active suspensions [22]. With orientations fixed, the near–field stresslet retains its $1/r$ form and generates a consistent position-based Hamiltonian, enabling a direct comparison with the far–field dynamics. Thus, in the unscreened regime ($\kappa r_{ij}\ll 1$), the stream function of dipole $j$ is given by Eq. (14), leading to the Hamiltonian
@@ V Collective dynamics of pusher and puller dipoles
We probe the dynamical consequences of the two Hamiltonians $H_{\rm far}$ and $H_{\rm near}$ by separately time-evolving ensembles of identical pushers and pullers ($\sigma_{i}=\pm 1$) with fixed, co–aligned orientations. In the far field, the flow is vorticity–free and orientations are naturally quenched, making $H_{\rm far}$ an exact position–based Hamiltonian. In the near field, where vorticity is finite, a Hamiltonian description applies only in the enforced quenched–orientation limit. Within these regimes, we contrast the collective organization produced by $H_{\rm far}$ and $H_{\rm near}$.
@@ V.1 Simulation details
The collective dynamics of active dipoles in an incompressible membrane depend crucially on whether interactions are mediated by the screened $H_{\rm far}$ or by the unscreened $H_{\rm{}_{n}ear}$. Figure 2 summarizes the behaviour for ensembles of $N=15$ pushers (top row) and pullers (bottom row).
@@
In contrast, evolution under the unscreened near–field Hamiltonian $H_{\rm near}$ [Figs. 2(c,d,g,h)] exhibits a completely different trend. The $1/r$ stresslet decay is slower and we find that the ensemble does not collapse. Instead, particles spread into extended configurations and $\langle d_{ij}\rangle(t)$ grows monotonically. In this near–zone regime, the self–aggregation observed in the far field is vastly suppressed: although individual pairs can transiently attract or repel, they fail to form many–body stable compact clusters .
The similarity between the pusher and puller panels highlights a central result of this work: the nature of the hydrodynamic kernel, rather than the motor type alone, determines whether active dipoles aggregate or disperse. The far–field screened flow is irrotational and generates an attractive effective Hamiltonian, while the near–field unscreened flow disfavors collective collapse. For completeness, Fig. 3 shows a representative far–zone evolution with arbitrary orientations, illustrating that the same aggregation mechanism persists beyond the co–aligned limit. Representative far- and near-zone trajectories of co-aligned pusher and puller clusters are provided in the Supplementary Material (Figs. S1–S2). In all these examples, a soft harmonic repulsion is added to regularize close encounters and suppress particle overlap.
@@ VI Summary and outlook
We have provided a concrete realization of a minimal hydrodynamic Hamiltonian framework for active force dipoles confined to an incompressible supported membrane. Starting from the Brinkman–regularized Stokes equation, we derived the screened incompressible Green’s tensor, identified its logarithmic near–field and algebraically screened far–field limits, and constructed the associated dipolar flows and stream functions. A key outcome of this analysis is the emergence of two distinct Hamiltonian constructions. In the far field, the screened flow is strictly vorticity–free, so dipole orientations remain naturally quenched; membrane incompressibility then yields an exact position–based Hamiltonian $H_{\rm far}$. In contrast, the near field possesses finite vorticity, precluding an orientation–dynamics–free description; a Hamiltonian formulation applies only in the enforced quenched–orientation limit, defining the effective Hamiltonian $H_{\rm near}$. Within these regimes, we contrast the collective organization generated by the two hydrodynamic Hamiltonians.
@@
This comparison reveals a striking dichotomy. For identical pushers (and similarly for pullers), the far–field Hamiltonian $H_{\rm far}$ produces a strong self–binding instability, driving rapid collapse into compact clusters and a sharp decay of the mean pair separation. The unscreened near–field Hamiltonian $H_{\rm near}$ behaves qualitatively differently: it induces strong local rearrangements but no long–lived aggregates, and the ensemble disperses monotonically in time, rendering near–zone dynamics intrinsically non–aggregating. This contrast persists across motor types and initial orientations, showing that it is the structure of the hydrodynamic Hamiltonians—not merely the dipole sign—that controls collective organization. Moreover, our work establishes the important roles of membrane vorticity and incompressibility that leads to the emergence of such Hamiltonian descriptions. These results offer a framework for interpreting more complex membrane rheologies. Allowing membrane compressibility opens a longitudinal mobility channel that breaks the Hamiltonian structure and can enhance aggregation for pullers [24] while producing mixed responses for pushers. Introducing odd (Hall) viscosity adds antisymmetric stresses that drive chiral transport and circulating cluster states. An important open question is the extent to which such effects can be incorporated within generalized Hamiltonian frameworks, [24]. Coupling the membrane to curvature, shape fluctuations, or biochemical activity further broadens the phenomenology, enabling a range of richer and biologically relevant collective behaviors, some of which may admit generalized Hamiltonian descriptions based on symmetry considerations. More broadly, incompressible supported membranes provide a minimal yet dynamically rich example in which explicit Hamiltonian descriptions of active particles emerge under suitable conditions. This perspective motivates experimental tests in reconstituted active films, where screening lengths and membrane viscosities can be precisely tuned, as well as theoretical extensions of Hamiltonian and symmetry-based approaches to active matter at interfaces.
@@ Acknowledgements.
The numerical implementations used in this study are available from the authors upon reasonable request.
Appendix A Supplementary Figures
(a) Pushers (far zone)
(b) Pullers (far zone)
(c) Pushers (near zone)
(d) Pullers (near zone)
Figure 4: Far- and near-zone dynamics of co-aligned clusters of twelve active particles in an incompressible membrane. Columns correspond to pushers (left) and pullers (right). Particles are initialized at random locations within a disc. Initial positions are shown in green (pushers) and yellow (pullers), final positions in red and orange, respectively, with trajectories shown in grey. A soft harmonic repulsion is included to prevent particle overlap.